### v1

# Also at ] The European Space Agency Centre for Earth Observation, Frascati, Italy

Abstract
Recent minimalist modeling has demonstrated that overdamped polar chiral active matter can support an emergent, inviscid Euler turbulence, despite the system’s strictly dissipative microscopic nature. In this letter, we establish the statistical mechanical foundation for the emergent inertial regime by deriving a formal isomorphism between the model’s agent dynamics and the overdamped Langevin equation for disordered Josephson junctions. We identify the trapped agent state as a macroscopic superconducting phase governed by the Adler equation. The validity of this mapping is confirmed analytically by a disorder-broadened Adler-Ohmic crossover in the system’s slip velocity, corresponding to the saddle-node bifurcation of phase-locking systems. These results define the new minimal chiral flocking model as a motile, disordered Josephson array, bridging active turbulence and quantum superconductivity.
preprint: APS/123-QED
I Introduction
We consider a recent, minimalist active matter turbulence model that is characterized by three simple rules, namely that its agents are (1) self-propelled, (2) polar, or flocking, favoring alignment, and (3) chiral, with intrinsic rotation [1, 2]. The model is unique in that it collapses these three properties into a single degree of freedom, phase, $\phi_{i}$,
$\dot{\mathbf{r}}_{i}=v_{0}\mathbf{\hat{n}}(\phi_{i}),$ (1)
where $\mathbf{r}_{i}$ is an agent’s position, $v_{0}$ its constant swim speed, and $\mathbf{\hat{n}}(\phi_{i})=(\cos\phi_{i},\;\sin\phi_{i})$ is the polar unit vector for internal phase $\phi_{i}\in[0,2\pi)$. The time evolution in phase is given by,
$\dot{\phi}_{i}=\omega_{i}+a_{0}R\sin(\Psi-\phi_{i})+\eta_{i}(t),$ (2)
where $\omega_{i}$ is an agent’s intrinsic chirality, or frustration, $\eta_{i}(t)$ is a Gaussian white noise term, and $a_{0}R\sin(\Psi-\phi_{i})$ is a localized Kuramoto-Sakaguchi-coupling [3, 4]. With $R$ being the local order parameter, we note that
$Z(\mathbf{x},t)=R(\mathbf{x},t)e^{i\Psi(\mathbf{x},t)}=\\ =\int_{\mathcal{D}}G(|\mathbf{x}-\mathbf{x}^{\prime}|)\left[\sum_{j=1}^{N}\delta(\mathbf{x}^{\prime}-\mathbf{x}_{j}(t))e^{i\phi_{j}(t)}\right]d\mathbf{x}^{\prime},$ (3)
is the local complex order parameter field, expressed as the convolution of the microscopic agent distribution with a finite-range interaction kernel $G(|\mathbf{x}-\mathbf{x}^{\prime}|)$, representing attractive forces between agents that align. By defining a continuum field theory for Eqs. (2) and (3), Ref. [1] demonstrated that the ensemble of active agents is subject to a kinetic Turing instability, triggered by the interplay between $Z$ (activation) and $\omega_{i}$ (inhibition).
Of some importance, Refs. [1] and [2] both explicitly considered a renormalized fluid element in the model, the former by calculating the guiding centre location of each agent’s oscillatory motion, and the latter by the implementation of a low-pass filter scaled by the interaction correlation timescale. The papers motivated this step in terms of the Vlasov-Fokker-Planck equation [1] and shallow water theory [2], and the latter rigorously demonstrated that Eqs. (2) and (3) thereby lead to inviscid Euler turbulence in the form of an Onsager dipole.
However, an important question remains. Our active agent ensemble is demonstrably overdamped and dissipative, yet the model exhibits an inverse, conservative energy cascade. How does a system that should fundamentally decay instead pump energy towards the largest scales? The resolution lies in considering the discrete distribution of agents, without a renormalized fluid element transformation, explicitly as a spatially extended, disordered Josephson junction array. The depinning transition between ’trapped’ and ’running’ states then generates a macroscopic synchronization stiffness, providing an effective inertia.
II Theoretical Foundation
Josephson junctions are essentially ensembles of quantum oscillators with a non-zero dispersion $\Delta\omega$ in their natural frequencies $\omega_{i}$, unavoidable due to fabrication heterogeneity [5, 6]. Josephson junctions synchronize (phase-lock) via a collective (global) coupling $K$ if the coupling strength is able to overcome the ensemble’s fabrication heterogeneity (frequency dispersion) $\Delta\omega$. Crucially, the coupling $K$ is naturally described with the Kuramoto-model [7], readily providing a theoretical foundation to understand the thermodynamics of our minimal model, which features a long-range, locally mediated Kuramoto interaction (Eqs. 2, 3).
In this letter, we shall consider Eq. (2) as the overdamped Langevin equation, following Ref. [8], which leads to the tilted washboard potential [6]. Given that our model’s agents are essentially massless, we shall derive a formal mapping to the resistively shunted junction model in the overdamped limit [9, 10].
We begin by writing Eq. (2) as,
$\dot{\phi}_{i}=-\frac{\partial V_{\text{eff}}}{\partial\phi_{i}}+\eta_{i}(t),$ (4)
where,
$V_{\text{eff}}(\phi_{i})=-a_{0}R\cos(\Psi-\phi_{i})-\omega_{i}\phi_{i},$ (5)
is the classic tilted washboard potential characterized by two competing energy scales, the potential wells $a_{0}R$ (which synchronize agents), and the tilt $\omega_{i}$ (which de-synchronize agents). From Eq. (5) we observe that agents can be ‘trapped’ when $a_{0}R>\omega_{i}$, ‘running’ when $a_{0}R<\omega_{i}$, and ‘marginal’ when $a_{0}R\approx\omega_{i}$. $\eta_{i}(t)$ (as well as the intrinsic $\omega_{i}$) act as sub-gap leakage [11], the background entropy production caused by noise and disorder. These terms drive an irreducible background entropy production [12], ensuring that the emergent inertial fluid retains a finite ’thermal viscosity’ even deep within the phase-locked regime.
To characterize the thermodynamics of Ref. [2]’s fluid-theoretically derived “shock cycle” between the active, dissipative bath and an ordered, inertial state, we shall derive an agent’s phase velocity in the co-moving Lagrangian frame of the local order parameter. We start by decomposing an agent’s phase velocity,
$\phi_{i}(t)=\Psi_{i}(t)+\delta\phi_{i}(t),$ (6)
where $\Psi_{i}(t)=\Psi(\mathbf{r}_{i}(t),t)$ is the local phase at agent $i$’s location, representing the collective phase evolution, and $\delta\phi_{i}(t)$ represents an agent’s relative phase phase eovlution. Taking the time derivative of Eq. (6) yields,
$\dot{\phi_{i}}=\dot{\Psi}+\dot{\delta\phi},$ (7)
from which we define,
$v_{\text{slip}}\equiv|\dot{\delta\phi_{i}}|=|\dot{\phi_{i}}-\dot{\Psi_{i}}|,$ (8)
as an agent’s phase slip velocity. Here, we identify $\dot{\phi_{i}}$ as an agent’s deterministic phase velocity (Eqs. 2, 5) and $\dot{\Psi_{i}}$ as the phase velocity of the potential well seen by that agent, meaning that $v_{\text{slip}}$ represents an agent’s phase speed relative to a potential well. By quantifying $v_{\text{slip}}$ (Eq. 8) we can therefore measure whether an agent is being thermalized into the active bath, or whether it is contributing to an inertial Euler flow in $R^{2}$ [2].
This realization allows us to pinpoint the mathematical isomorphism that explains why the system bifurcates into a hydrodynamic (trapped) state and an active bath (running) state. We use Eq. (2) deterministically to rewrite Eq. (7) in form of,
$\dot{\delta\phi_{i}}=\omega_{i}-\dot{\Psi_{i}}-a_{0}R\sin\delta\phi,$ (9)
which is, in exact terms, the Adler equation [13, 14] for a dirty, or disordered, Josephson array [7], with individual junctions driven by an effective current $I_{\text{eff}}$ [15],
$I_{\text{eff}}=\omega_{i}-\dot{\Psi_{i}}.$ (10)
Here, we identify $\Delta\omega$,
$\langle\omega_{i}-\dot{\Psi_{i}}\rangle=\Delta\omega,$ (11)
the dispersion in the agents’ intrinsic frequency, corresponding to inertial fluid’s topography in the shallow water isomorphism derived in Ref. [2]. Consistent this inference, the effective current in Eq. (10) is a driving force that tries to break local synchronization. Setting Eq. (9) to zero and solving for $\delta\phi$ yields the critical current,
$I_{c}=a_{0}R,$ (12)
which can likewise be derived from considering the potential well depths in the tilted washboard potential (Eq. 5). When $|I_{\text{eff}}|<I_{c}$, the local coupling force $a_{0}R$ is strong enough to trap the agent, which subsequently locks onto a stable fixed angle and swims with the local group. Conversely, when $|I_{\text{eff}}|>I_{c}$, the agent’s intrinsic frustration exceeds the local maximum coupling force. The agent runs or slips relative to the group, with the phase continuously rotating.
Figure 1: A simulated ensemble of polar chiral active matter turbulence. Panels a–d) show end-state snapshots of the entire simulation space, with agent point-cloud (panel a), topological charge (panel b), flow-lines of circulation in order $R^{2}$ (panel c), and agent slip velocity $v_{\text{slip}}$ (panel d), with respective colorscales where applicable. Panel e) shows defect count as a function of normalized simulation time (left $y$-axis) with a $t^{-0.75}$ powerlaw fit shown with a green, dashed line, and the percentage of trapped agents shown in red (right $y$-axis, inverted), using the heuristic definition of a trapped agent having a slip velocity below 0.05 rad/s. Panel f) shows a cross-correlation analysis of the latter, for normalized times >4.3>4.3, using the $t^{-0.75}$ powerlaw to detrend the defect count timeseries. Estimated confidence intervals for the cross-correlation are shown with dashed, black lines. Panels g–j) show inset $v_{\text{slip}}$-snapshots in rapid succession, highlighting ongoing shock merger, agent thermalization, and subsequent re-organization.
In equivalent terms, for Josephson junctions, when $I_{\text{eff}}<I_{c}$ an oscillator famously locks onto a constant value, with $I_{c}$ there being a material quantity dependent on the semiconductor in question. However, whereas the critical current of a Josephson junction dependends on material qualities, $I_{c}=a_{0}R$ is in the present model entirely dynamic and depends on local synchronization. Still, we observe complete isomorphism between Eq. (9) and the Adler equation for a Josephson junction with in-series resistance [15]. When $I_{\text{eff}}>I_{c}$ the Josephson junction cannot hold onto the phase, acquires a voltage, and subsequently produce entropy through dissipative currents. Equivalently, in our model, when an agent runs or slips relative to the group, it acquires the voltage,
$V_{\text{eff}}=|\dot{\phi}_{i}-\dot{\Psi}_{i}|=v_{\text{slip}}.$ (13)
The resistive motion of the agent then thermalizes energy into the active bath.
We can thus formalize Ref. [2]’s description of the model’s central thermodynamic process; when agents start to slip, $R$ drops, lowering the local critical current $I_{c}$. This causes neighbours to slip, forcing a cascade in energy, resulting in a locally disorganized state ($R\approx 0$). However, this state is linearly unstable to the kinetic Turing instability [1]. Once agents have thermalized, they re-organize. In the tilted washboard formalism, as $R$ grows, the potential wells deepen, drawing more agents in. Their intrinsic rotation $\omega_{i}$ is now synchronized by spatial segregation, producing a distinct topography for the inertial flow [2].
Following Ref. [15]’s demonstration of how individual Josephson oscillators respond to external forcing, we calculate the time required for a $2\pi$ phase slip,
$T_{\text{slip}}=\int_{0}^{2\pi};\frac{d\delta\phi}{\dot{\delta\phi}}=\frac{2\pi}{\sqrt{I_{\text{eff}}^{2}-I_{c}^{2}}},$ (14)
where we inserted our Adler equation into the integral (Eq. 9). Consequently, the instantaneous slip velocity is
$v_{\text{slip}}=\frac{2\pi}{T_{\text{slip}}}=\sqrt{I_{\text{eff}}^{2}-I_{c}^{2}}.$ (15)
Examining the asymptotic limits of Eq. (15) reveals two distinct dynamical regimes in the deterministic limit (see also Ref. [16]):
$v_{\text{det}}=\begin{cases}\sqrt{I_{\text{eff}}^{2}-I_{c}^{2}}\sim(I_{\text{eff}}-I_{c})^{1/2}&\text{ for }I_{\text{eff}}\geq I_{c}\\ \\ 0&\text{ for }I_{\text{eff}}<I_{c}\end{cases}$ (16)
In equivalent terms, we follow Ref. [15]’s formalism for the ensemble average, $\langle v_{\text{slip}}\rangle$, obtained by integrating Eq. (15) over the probability distribution $P(I_{\text{eff}})$. This accounts for noise as a fluctuating tilt that statistically drives the system across the critical threshold:
$\langle v_{\text{slip}}\rangle=\int_{|I_{\text{eff}}|>I_{c}}P(I_{\text{eff}})\sqrt{I_{\text{eff}}^{2}-I_{c}^{2}},dI_{\text{eff}}.$ (17)
Eq. (17) is expected to conform to the specific ’soft knee’ geometry identified in Ref. [15] and stands in direct correspondence with Ref. [7], which established that the macroscopic resistance of a disordered Josephson array scales with the fraction of running junctions, demarcating the phase transition from a ‘superconducting glass’ (pinned, or trapped, phases) to a resistive metal (running phases).
In what follows, we demonstrate that our minimal active matter model effectively realizes a motile Josephson array. By embedding the discrete oscillator dynamics of Ref. [15] into intrinsically motile (self-propelled) agents, we introduce a transport mechanism absent in standard static arrays [7]. In our model, dissipative excitations are physically advected by the running agents themselves. This active transport expels disorder from the ordered bulk and concentrates it into defect cores, segregating the system into a dual-phase fluid. The segregation mimics the thermodynamic state preparation observed in driven optical systems [17], where the interplay of drive and dissipation stabilizes distinct macroscopic phases (see also Ref. [18]). Specifically, we observe (1) an ideal, inviscid Euler fluid of trapped agents with inertial mass proportional to order squared [2], and (2) a dissipative resistive fluid of running agents. This resistive phase acts as the thermodynamic sink that ultimately sustains the cycle, permitting the regeneration of order via a kinetic Turing instability [1].
III Results
We have implemented the model (Eqs. 2, 3) in a bounded periodic box with $N=50,000$ agents, akin to Refs. [1, 2]. Figure 1 presents snapshots of the fully formed Onsager dipole, with panels a–c) showing agent point cloud, topological charge, and inertial flowlines, respectively. Panel d) shows the slip velocity $v_{slip}$ (Eq. 8), for the entire simulation space.
Figure 1g–j), showing four prior $v_{slip}$-snapshots of a limited portion of the simulation space, offer an instructional explanation of the model’s core thermodynamic mechanism. An ordered ($R\approx 1$) region shatters into phase slips (active turbulence), generating entropy; this is the shock merger mechanism identified by Ref. [2]. The agents inside this now-disordered patch are then quickly reorganized into the inertial fluid by the kinetic Turing instability [1]. It is this process that acts as a thermodynamic state preparation cycle [17], effectively ’pumping’ the system from the high-entropy running state back into the low-entropy trapped state, sustaining the Onsager dipole.
Figure 1e, f) demonstrate the foregoing dynamic in rigorous terms. Panel e) shows defect count going down with a characteristic scaling of $t^{-0.75}$ [19, 20], identified by Ref. [2] to be caused by shock dynamics. Predictably, the percentage of trapped agents increase from the initial chaos [18] (red line in Figure 1e, right $y$-axis, inverted). However, subtracting the $t^{-0.75}$ scaling from the defect count reveal that the fluctuations are moderately correlated (Figure 1f, cross-correlation). The apparent discrepancy is resolved by recognizing that, as topological defects annihilate, the kinetic energy stored in the vortex cores is released into the bulk (Figure 1g–j). A reduction in topological defects is therefore immediately followed by a transient decrease in the global trapped agent fraction due to noise-induced delocking [21, 15]. This is consistent with a latent heat release during a dynamical phase transition [7], generating a thermal spike that exponentially and momentarily increases the probability of phase slips.
Figure 2: The result of 72 simulations systematically varying the exponent in the $\omega_{i}$ distribution, yielding a systematically varied $\Delta\omega$, the dispersion in intrinsic agent frustration $\omega_{i}$ (inset). The main plot shows observed ensemble average of the slip velocity $\langle v_{\text{slip}}\rangle$ (voltage) against dispersion $\Delta\omega$ (current), in a log-log “I-V” plot. $\langle v_{\text{slip}}\rangle$ is calculated as the mean value of $v_{\text{slip}}$ for all NN agents, collected during four snapshots of the second half of each simulation. A solid black line shows a $4^{\text{th}}$ order Hermite polynomial fit, with a dashed-black line indicating the knee identified by non-linear least squares minimization [22]. Dashed-blue line shows a $2^{\text{nd}}$ order (piece-wise linear) Hermite polynomial fit with slopes (exponent $\beta$) of 0 and 1, with a blue-dotted line indicating the theoretical mean-field critical threshold dispersion, $\Delta\omega_{c}\approx 9\,\text{s}^{-1}$. The observable onset of slip velocity occurs significantly earlier (black-dashed line, at $\Delta\omega\approx 2\,\text{s}^{-1}$), indicating substantial sub-gap leakage current driven by the intrinsic disorder of the array.
We formalize the foregoing description by systematically varying $\langle I_{\text{eff}}\rangle=\Delta\omega$ (Eq. 11) in successive simulations, essentially implementing a Monte-Carlo simulation of the current-voltage (I-V) curve for the motile Josephson array under consideration. Here, we note that the dispersion $\Delta\omega$ is entirely decided by the exponent $n$ of the powerlaw distribution in $\omega_{i}$; $P(\omega)\sim\omega^{-n}$ [1, 2]. We therefore systematically vary $n$ from $0.5$ to $4$. Figure 2 shows the result in a scatterplot of the ensemble-averaged slip velocity $\langle v_{\text{slip}}\rangle$ against $\Delta\omega$. To characterize the macroscopic response function without imposing the rigid constraints of a single-parameter junction model, we fit $\langle v_{\text{slip}}\rangle$ using a shape-preserving $4^{\text{th}}$-order Hermite polynomial [22] (solid black line). This phenomenological approach accommodates the inhomogeneous broadening caused by the structural disorder of the agent array [15].
Crucially, the solid-black fit in Figure 2 reveals a continuous evolution of the local scaling exponent. At very low dispersion ($\Delta\omega\ll\Delta\omega_{c}$), the response is dominated by a noise floor ($\beta=0$), gradually rising towards $\Delta\omega_{c}\approx 9$ s-1. The $4^{\text{th}}$-order Hermite polynomial fit identifies a knee at $\Delta\omega\approx 2$ s-1, and the relatively large spread between these values is consistent with sub-gap leakage currents smoothing out the $\beta=0.5$ transition [15]. As the dispersion exceeds the critical threshold, $\Delta\omega>\Delta\omega_{c}$, the system quickly and strictly adheres to a linear regime, asymptotically recovering the Adler-Ohmic limit ($\beta=1$).
IV Summary & Discussion
In this letter, we have derived a strict isomorphism between a minimalist active matter model to disordered Josephson junctions. In so doing, we have resolved the origin of emergent inertial flows in overdamped chiral flocks [1, 2]. We demonstrate that the system is governed by a tilted washboard potential, where a thermodynamic transition separates resistive ‘running’ agents from the ‘trapped’ population that sustain a self-organizing inertial cascade. This establishes the minimalist active matter model as a distributed, motile Josephson array.
The classification of the minimal model as a motile Josephson array is empirically demanded by the system’s macroscopic response function (Figure 2), which traverses the distinct Adler-Ohmic crossover predicted by the resistively shunted junction formalism [16, 7] with soft knee-geometry identified in Ref. [15], the topological fingerprint of the Adler equation’s saddle-node bifurcation, definitively distinguishing the mechanism from standard viscous fluid transitions and justifying the identification of a tilted washboard potential [8]. The subsequent recovery of the linear Ohmic limit ($\beta=1$) at high dispersion ($\Delta\omega$) confirms that Ref. [2]’s discovery of the system’s ’inertial’ phenomenology is rigorously governed by the non-linear phase topology of disordered oscillator arrays.
The formal isomorphism to the standard resistively shunted junction model strictly relies on the polar symmetry of the interaction (flocking) [23]. nematic active matter (half-integer symmetry) would imply a $\pi$-periodic potential [24], mapping to distinct universality classes (e.g., XY2 models [25]) that do not support the integer vortex topology of the Onsager dipole. In contrast, the $2\pi$-periodicity of the polar alignment interaction $\sin(\phi_{i}-\Psi)$ must map directly to the sinusoidal current-phase relation of a superconducting tunnel junction [5, 6].
By bridging chiral active matter turbulence and superconductivity we have identified phase synchronization as a microscopic engine of macroscopic order, establishing the motile Josephson array as a new universality class of driven chiral flocks.
Acknowledgements
This work is supported by the European Space Agency’s Living Planet Grant No. 1000012348. The author is grateful to O. Nestande, D. Knudsen, PT. Jayachandran, and K. Douch for stimulating discussions. Google’s Gemini 3.0 Pro has been used for mathematical formalism and coding assistance in matlab.